OpenStudy (anitaasks):

How do I find the limit as x approaches negative infinity of x^3e^x? I know that I am supposed to use L'Hopital's rule, but I'm not sure how

Miracrown (miracrown):

\[\lim_{x \rightarrow - \infty} \space ^{x ^{3 \space e^x}} \]

Miracrown (miracrown):

Do you have any idea on how to approach this problem?

Miracrown (miracrown):

Youu are correct about using L Hospital rule However, to use L Hospital rule, we will convert our function into the fraction form

Miracrown (miracrown):

\[= \lim_{x \rightarrow \infty} \space \frac{ x ^{3} }{ e ^{-x} }\]

Miracrown (miracrown):

if I bring e^x to the bottom, its exponent will change sign right?

Miracrown (miracrown):

Now let's take derivative of both top and bottom, would you like to give it a try?

OpenStudy (anitaasks):

Yeah I tried that, and I got 3x^2/-e^-x

Miracrown (miracrown):

Excellent!

OpenStudy (dan815):

http://prntscr.com/6e3iq3

Miracrown (miracrown):

\[\lim_{x \rightarrow \infty} \space \frac{ 3x ^{2} }{ -e ^{-x}}\]

Miracrown (miracrown):

If we evaluate this limit when x approaches negative infinity, do you have any idea on what this fraction approaches?

OpenStudy (anitaasks):

Infinity over -infinity?

Miracrown (miracrown):

That's right :)

Miracrown (miracrown):

So it is still indeterminate form right?

OpenStudy (anitaasks):

Yeah so I did the rule twice more and got 6/-e^-x, but that didn't really work

Miracrown (miracrown):

Great! It looks good to me

Miracrown (miracrown):

\[= \lim_{x \rightarrow \infty} \space \frac{ 6 }{ -e ^{-x} }\]

Miracrown (miracrown):

Now we don't have indefinite form anymore because the top is fixed at 6 what will the bottom approach?

OpenStudy (anitaasks):

-infinity?

Miracrown (miracrown):

Exactly

Miracrown (miracrown):

if the bottom is huge and the top is small, then what does the whole fraction approach?

OpenStudy (anitaasks):

OHH I see it now! It's zero. I forgot about this property, haha. Thank you so much!

Miracrown (miracrown):

Haha, no worries. =]

OpenStudy (dan815):

uhh

OpenStudy (dan815):

why did u say the bottom is going to infinity?

OpenStudy (dan815):

http://prntscr.com/6e3mci

OpenStudy (dan815):

there is some error in the solution method

OpenStudy (dan815):

i think what u want to write is this

OpenStudy (dan815):